# Emergence of Vorticity and Viscous Stress in Finite Scale Quantum Hydrodynamics

### v1

@@ IV Example Flow Field: A Line Vortex
$\omega(r)=\omega_{0}\left[1-\frac{1}{2}\frac{r^{2}\sigma^{2}}{\ell^{2}\left(\sigma^{2}+\ell^{2}\right)}\right].$ (35)
In the opposite limit, $r>>\alpha$ we note that $\frac{r}{\alpha}e^{-\frac{r^{2}}{\alpha^{2}}\sin^{2}\phi}\approx\sqrt{\pi}\delta\left(\sin\phi\right)$, so that:
$\lim_{r\rightarrow\infty}\omega(r)$ $=\omega_{0}\left\{1-\frac{1}{2}\int_{0}^{2\pi}d\phi\cos\phi\right$. (36)
$\left.\times\left[\delta\left(\phi\right)+\delta\left(\phi-\pi\right)\right]\lim_{r\rightarrow\infty}\text{erf}\left(\frac{r\cos\phi}{\alpha}\right)\right\},$
$=0.$
This demonstrates that the finite scale coarse graining procedure can indeed give rise to finite vorticity, even though the underlying microscopic flow is irrotational. This vorticity is peaked at the vortex core and decreases to zero at large radii compared to $\alpha=\ell\sqrt{2\left(\sigma^{2}+\ell^{2}\right)}/\sigma$, see Figure 1. Furthermore, we note that in the limit of $\ell\rightarrow 0$ the peak at the vortex core diverges while the value at any finite radius away from the core vanishes, in agreement with the singular behavior of the microscopic flow field. In this sense, we can think of the coarse graining procedure as smoothing out the singular behavior of the line vortex and distributing the circulation into a region with finite vorticity.
@@ V Discussion and Conclusions
In this paper we have derived a set of hydrodynamic equations for spatially coarse-grained quantum systems. The coarse-graining procedure we employed represents a generalization of the finite scale theory which has been previously applied in the study of classical fluids. As in the case of classical fluids, we find that the appropriate velocity field in the macroscopic fluid equations is a Favre averaged velocity field and that the coarse-graining gives rise to emergent viscous stresses in the momentum equations. We have also shown that this macroscopic velocity field can possess an emergent vorticity, even though the microscopic velocity is irrotational. The equation governing the evolution of the vorticity was also derived and shown to possess the same vortex-stretching term that is present in classical fluids. Furthermore, expanding the stress terms to leading order in small parameters, we found that the equation governing the vorticity evolution depends only on coarse-grained quantities and has no dependence on the parameters describing the microscopic quantum model, consistent with its emergence at finite scales.
The existence of finite vorticity was demonstrated explicitly in Sec.IV using the specific example of an irrotational line vortex. It was shown that the emergent vorticity in that example is proportional to the vortex strength and that the coarse-graining acts to smooth out the singular behavior of the underlying microscopic velocity field and distribute finite vorticity through the fluid around the vortex. It should be noted that, in that case, the small parameter associated with the coarse-graining length scale appeared in the denominator of the vorticity due to singular behavior in the microscopic theory. This implies that for microscopic fields with singularities, the finite scale expansion presented in this work may require modifications. However, while the small parameter expansion may fail for singular microscopic fields, the exact expressions presented in this work should still apply and the general conclusions about the emergence of vorticity and novel viscous stress terms in the hydrodynamic equations remain valid even for singular fields.

### v2

@@ IV Example Flow Field: A Line Vortex
$\boldsymbol{\omega}=\hat{z}\frac{\Gamma}{\pi}\frac{2\ell^{2}}{\left(2\ell^{2}+\tfrac{R^{2}\sigma^{2}}{\ell^{2}+\sigma^{2}}\right)^{2}}.$ (39)
At $R=0$, the vorticity is again $\boldsymbol{\omega}=\hat{z}\Gamma/2\pi\ell^{2}$, just as the $n=0$ case. However, for $R,\sigma>0$, and $\ell<<R,\sigma$, the vorticity becomes:
$\boldsymbol{\omega}=\hat{z}\frac{2\Gamma}{\pi}\frac{\ell^{2}}{R^{4}}+\mathcal{O}(\ell^{4}).$ (40)
Comparing this to the finite scale expansion in Eq. (24), we confirm that the leading-order contributions for this case are proportional to $\ell^{2}$. This offers a concrete example of a vorticity field for which the finite-scale equations in Sec. III should capture the leading-order physics in small $\ell$.
Figure 1: Plot of emergent vorticity as given by Eq. (33) as a function of radius from vortex center, $R$, normalized to the length scale $\sigma$. Vorticity is normalized to $\omega_{0}=\Gamma/2\pi\ell^{2}$, and $\ell/\sigma$ is chosen to be $1/10$ for illustrative purposes.
Figure 1 shows a comparison of the emergent vorticity field, computed using the general formula, Eq. (33), for $n=0,1,2,3$. Note that in all four cases, the vorticity is peaked at the vortex core and decreases to zero at large radii $R/\sigma$. Moreover, we see that the width of the peak becomes narrower as $n$ increases, and, correspondingly, the tail drops off more slowly. In this sense, we can think of the coarse graining procedure as smoothing out the singular behavior of the line vortex and distributing the circulation into a region with finite vorticity.
@@ V Discussion and Conclusions
In this paper we have derived a closed set of hydrodynamic equations for spatially coarse-grained quantum systems. The coarse-graining procedure we employed leads to an infinite hierarchy of filtered moments, equivalent to what has been discussed previously in the literature. This hierarchy is then closed using a generalization of finite scale theory which has been previously applied in the study of classical fluids. Consistent with previous results in classical fluids, as well as recent work on quantum fluidsTanogami (2021), we find that the appropriate velocity field in the macroscopic fluid equations is a Favre averaged velocity field and that the coarse-graining gives rise to emergent viscous stresses in the momentum equations. After applying the finite-scale closure, these viscous stresses have the appearance of an artificial viscosity term, similar to those used in computational fluid dynamics. We have also shown that this macroscopic velocity field can possess an emergent vorticity, even though the microscopic velocity is irrotational. The equation governing the evolution of the vorticity was also derived and shown to possess the same vortex-stretching term that is present in classical fluids. Furthermore, expanding the stress terms to leading order in small parameters, we found that the equation governing the vorticity evolution depends only on coarse-grained quantities and has no dependence on the parameters describing the microscopic quantum model, consistent with its emergence at finite scales.
The existence of finite vorticity was demonstrated explicitly in Sec.IV using the specific example of an irrotational line vortex. It was shown that the emergent vorticity in that example is proportional to the vortex strength and that the coarse-graining acts to smooth out the singular behavior of the underlying microscopic velocity field and distribute finite vorticity through the fluid around the vortex. Importantly, when the density field is chosen to vanish at the vortex core, as it should in real systems, the leading order terms for the emergent vorticity are quadratic in the coarse-graining length scale, exactly as they appear in the finite-scale theory presented in Sec. III. Therefore, we expect that the finite-scale closure used here should capture the vortex dynamics of real quantized vortices.
@@ Acknowledgements.
Appendix A Line Vortex Coarse-Graining
In this appendix the closed form expressions for a coarse-grained line vortex are derived. Assume a wavefunction:
$\psi_{0}=\frac{\left(\textbf{x}_{\perp}\cdot\textbf{x}_{\perp}\right)^{n/2}}{(2\pi)^{3/4}\sigma^{3/2}r_{0}^{n}}e^{-\frac{\textbf{x}\cdot\textbf{x}}{4\sigma^{2}}}e^{i\frac{m}{\hbar}\frac{\Gamma}{2\pi}\tan^{-1}\left(\frac{x_{2}}{x_{1}}\right)},$ (41)
where $r_{0}$ is chosen to normalize the wavefunction. For $n\in\mathbb{N}$:
$r_{0}=\sqrt{2}\sigma\left(n!\right)^{\frac{1}{2n}}.$ (42)
$\rho$ $=\frac{m\left(\textbf{x}_{\perp}\cdot\textbf{x}_{\perp}\right)^{n}}{(2\pi)^{3/2}2^{n}\sigma^{2n+3}n!}e^{-\frac{\textbf{x}\cdot\textbf{x}}{2\sigma^{2}}},$ (43)
u $=\frac{\Gamma}{2\pi}\frac{-x_{2}\hat{x}_{1}+x_{1}\hat{x}_{2}}{\textbf{x}_{\perp}\cdot\textbf{x}_{\perp}},$
where $\Gamma$ is the circulation around the line vortex.
For the following coarse-graining analysis, a Gaussian filter will be used, which we write as:
$\left\langle\rho\right\rangle$ $=\int d^{3}x^{\prime}f\left(\textbf{x}^{\prime}\right)\rho\left(\textbf{x}^{\prime}+\textbf{x}\right),$
$=\frac{m}{(2\pi)^{3}\ell^{3}2^{n}\sigma^{2n+3}n!}\int_{-\infty}^{\infty}dz^{\prime}e^{-\frac{z^{\prime 2}}{2\sigma^{2}}}e^{-\frac{z^{\prime 2}+z^{2}-2zz^{\prime}}{2\ell^{2}}}$
$\times\int d^{2}x^{\prime}_{\perp}\left(\textbf{x}^{\prime}_{\perp}\cdot\textbf{x}^{\prime}_{\perp}\right)^{n}e^{-\frac{\textbf{x}^{\prime}_{\perp}\cdot\textbf{x}^{\prime}_{\perp}}{2\sigma^{2}}}e^{-\frac{\left(\textbf{x}^{\prime}_{\perp}-\textbf{x}_{\perp}\right)\cdot\left(\textbf{x}^{\prime}_{\perp}-\textbf{x}_{\perp}\right)}{2\ell^{2}}},$
$=\frac{me^{-\frac{z^{2}}{2\ell^{2}}\left(1-\frac{\tilde{\sigma}^{2}}{\ell^{2}}\right)}\tilde{\sigma}\sqrt{2\pi}}{(2\pi)^{3}\ell^{3}2^{n}\sigma^{2n+3}n!}e^{-\frac{R^{2}}{2\ell^{2}}}$
$\times\int_{0}^{2\pi}d\phi\int_{0}^{\infty}dr\ r^{2n+1}e^{-\frac{r^{2}}{2\tilde{\sigma}^{2}}}e^{\frac{rR\cos\phi}{\ell^{2}}},$
where $\tilde{\sigma}=\frac{\sigma\ell}{\sqrt{\ell^{2}+\sigma^{2}}}$. Changing variables and using the definition of the modified Bessel functions $I_{n}$ we find:
$\left\langle\rho\right\rangle$ $=\frac{me^{-\frac{z^{2}}{2\ell^{2}}\left(1-\frac{\tilde{\sigma}^{2}}{\ell^{2}}\right)}\tilde{\sigma}\sqrt{2\pi}}{(2\pi)^{2}\ell^{3}2^{n}\sigma^{2n+3}n!}e^{-\frac{R^{2}}{2\ell^{2}}}\left(2\tilde{\sigma}^{2}\right)^{n+1}$ (45)
$\times\int_{0}^{\infty}dt\ t^{n}e^{-t}I_{0}\left(2\kappa\sqrt{t}\right),$
where we define $\kappa\equiv R\tilde{\sigma}/\sqrt{2}\ell^{2}$. Note that the integral on the right hand side can be written exactly in terms of special functionsGradshteyn and Ryzhik (2014):
$\int_{0}^{\infty}$ $x^{\mu-\frac{1}{2}}e^{-\alpha x}I_{2\nu}\left(2\beta\sqrt{x}\right)dx=\frac{\Gamma\left(\mu+\nu+\tfrac{1}{2}\right)}{\Gamma\left(2\nu+1\right)}$ (46)
$\times\frac{\beta^{2\nu}}{\alpha^{\mu+\nu+\tfrac{1}{2}}}M\left(\mu+\nu+\tfrac{1}{2},2\nu+1,\frac{\beta^{2}}{\alpha}\right),$
where $M$ is the confluent hypergeometric function, and $\text{Re}\left\{\mu+\nu+\tfrac{1}{2}\right\}>0$.
It is straightforward to see that the integral in Eq. (45) can be evaluated using Eq. (46) with the identification: $\mu=n+\tfrac{1}{2}$, $\alpha=1$, $\nu=0$, and $\beta=\kappa$. This yields:
$\left\langle\rho\right\rangle=e^{-\frac{R^{2}}{2\ell^{2}}}e^{-\frac{z^{2}}{2\ell^{2}}\left(1-\frac{\tilde{\sigma}^{2}}{\ell^{2}}\right)}\frac{m\left(2\tilde{\sigma}^{2}\right)^{n+3/2}}{(4\pi)^{3/2}\ell^{3}2^{n}\sigma^{2n+3}}M\left(n+1,1,\kappa^{2}\right).$ (47)
Turning our attention to the Favre-averaged velocity field, we write the numerator as:
$\left\langle\rho\textbf{u}\right\rangle$ $=\frac{m\Gamma e^{-\frac{z^{2}}{2\ell^{2}}\left(1-\frac{\tilde{\sigma}^{2}}{\ell^{2}}\right)}\tilde{\sigma}\sqrt{2\pi}}{(2\pi)^{4}\ell^{3}\sigma^{3}r_{0}^{2n}}\int d^{2}x_{\perp}^{\prime}\left(\textbf{x}^{\prime}_{\perp}\cdot\textbf{x}^{\prime}_{\perp}\right)^{n-1}$
[…]
$=\frac{m\Gamma e^{-\frac{z^{2}}{2\ell^{2}}\left(1-\frac{\tilde{\sigma}^{2}}{\ell^{2}}\right)}\tilde{\sigma}\sqrt{2\pi}}{(2\pi)^{4}\ell^{3}\sigma^{3}r_{0}^{2n}}$
$\times\int_{0}^{2\pi}d\phi\int_{0}^{\infty}dr\ r^{2n}e^{-\frac{r^{2}}{2\sigma^{2}}}e^{-\frac{r^{2}+R^{2}-2rR\cos\phi}{2\ell^{2}}}$
$\times\left[-\sin\left(\phi+\theta\right)\hat{x}+\cos\left(\phi+\theta\right)\hat{y}\right$],
$=\frac{m\Gamma e^{-\frac{z^{2}}{2\ell^{2}}\left(1-\frac{\tilde{\sigma}^{2}}{\ell^{2}}\right)}\tilde{\sigma}\sqrt{2\pi}}{(2\pi)^{4}\ell^{3}\sigma^{3}r_{0}^{2n}}e^{-\frac{R^{2}}{2\ell^{2}}}$
$\times\int_{0}^{2\pi}d\phi\int_{0}^{\infty}dr\ r^{2n}e^{-\frac{r^{2}}{2\tilde{\sigma}^{2}}+\frac{rR}{\ell^{2}}\cos\phi}$
$\times\left[-\sin\phi\hat{r}+\cos\phi\hat{\theta}\right$],
where $\theta$ is defined by $\textbf{x}_{\perp}\cdot\hat{x}=R\cos\theta$. Note that the term proportional to $\hat{r}$ vanishes after performing the $\phi$-integral. Dividing this result by $\left\langle\rho\right\rangle$, and working through the algebra, we find:
$\left\langle\left\langle\textbf{u}\right\rangle\right\rangle$ $=\frac{\Gamma}{2\pi\tilde{\sigma}n!M\left(n+1,1,\frac{R^{2}\tilde{\sigma}^{2}}{2\ell^{4}}\right)}\frac{1}{\sqrt{2}}$
$\times\int_{0}^{\infty}dt\ t^{n-\tfrac{1}{2}}e^{-t}I_{1}\left(2\kappa\sqrt{t}\right)\ \hat{\theta}$.
The integral on the right hand side can be evaluated using the identity in Eq. (46), and identifying: $\mu=n$, $\alpha=1$, $\nu=\tfrac{1}{2}$, $\beta=\kappa$, leading to:
$\left\langle\left\langle\textbf{u}\right\rangle\right\rangle=\frac{R\ \Gamma}{4\pi\ell^{2}}\frac{M\left(n+1,2,\kappa^{2}\right)}{M\left(n+1,1,\kappa^{2}\right)}\ \hat{\theta}$. (48)
Using this expression, we can write the vorticity as:
$\boldsymbol{\omega}$ $=\boldsymbol{\nabla}\times\left\langle\left\langle\textbf{u}\right\rangle\right\rangle=\frac{1}{R}\hat{z}\frac{\partial}{\partial R}\left(R\left\langle\left\langle u_{\theta}\right\rangle\right\rangle\right),$
$=\hat{z}\frac{\Gamma}{4\pi\ell^{2}}\left[\frac{M\left(n+1,2,\kappa^{2}\right)}{M\left(n+1,1,\kappa^{2}\right)}+\frac{\partial}{\partial R}R\frac{M\left(n+1,2,\kappa^{2}\right)}{M\left(n+1,1,\kappa^{2}\right)}\right].$
This can be expressed exactly in terms of just $M\left(n+1,1,\kappa^{2}\right)$ and $M\left(n+1,2,\kappa^{2}\right)$ using the general relation from the NIST Digital Library of Mathematical Functions (13.3.15)DLMF :
$\frac{dM(a,b,z)}{dz}=\frac{a}{b}M(a+1,b+1,z),$ (49)
together with the contiguous relations for M(a,b,z)M(a,b,z)DLMF . While the analysis contained in this appendix only makes use of three of the contiguous relations, all six are included for completeness:
$(b-a)M(a-1,b,z)+(2a-b+z)M(a,b,z)-aM(a+1,b,z)$ $=0,\ \ (i)$ (50)
$b(b-1)M(a,b-1,z)+b(1-b-z)M(a,b,z)+z(b-a)M(a,b+1,z)$ $=0,\ \ (ii)$
$(a-b+1)M(a,b,z)-aM(a+1,b,z)+(b-1)M(a,b-1,z)$ $=0,\ \ (iii)$
$bM(a,b,z)-bM(a-1,b,z)-zM(a,b+1,z)$ $=0,\ \ (iv)$
$b(a+z)M(a,b,z)+z(a-b)M(a,b+1,z)-abM(a+1,b,z)$ $=0,\ \ (v)$
$(a-1+z)M(a,b,z)+(b-a)M(a-1,b,z)+(1-b)M(a,b-1,z)$ $=0.\,\ \ (vi)$
These six relations can be used to connect $M(a,b,z)$ to the four functions: $M(a\pm 1,b,z)$ and $M(a,b\pm 1,z)$. For the purposes of this appendix, we combine the $(vi)$ with $(v)$ to arrive at:
$M(a+1,b+1,z)=\frac{a-b}{a}M(a,b+1,z)+\frac{b}{a}M(a,b,z),$ (51)
and also:
$M(a+1,b+1,z)=\frac{b}{z}\left[M(a+1,b,z)-M(a,b,z)\right].$ (52)
Combining $(iii)$ with Eq. (52) we find:
$M(a+1,b+1,z)=\frac{b(b-1)}{az}\left[M(a,b-1,z)-M(a,b,z)\right].$ (53)
Using Eq. (49), together with Eqs. (51) and (53), it is straightforward to show that:
$\boldsymbol{\omega}=\hat{z}\frac{\Gamma}{2\pi\ell^{2}}\left\{1-\kappa^{2}\frac{M\left(n+1,2,\kappa^{2}\right)}{M\left(n+1,1,\kappa^{2}\right)}\left[1+n\frac{M\left(n+1,2,\kappa^{2}\right)}{M\left(n+1,1,\kappa^{2}\right)}\right]\right\}.$ (54)
The general form for the coarse-grained vorticity, Eq. (54), can be used to evaluate the emergent vorticity for any vortex line described by the general form in Eq. (41), where $\rho\sim R^{2n}$ near the vortex core. Notice that the general forms for the coarse-grained density, velocity, and vorticity, Eqs. (47), (48), and (54), are written in terms of just $M(n+1,1,\kappa^{2})$ and $M(n+1,2,\kappa^{2})$.
Let us now consider two special cases, for which the confluent hypergeometric functions reduce to elementary functions.
A.1 Special Case: n=0
The case of $n=0$ corresponds to the density field:
$\rho=\frac{m}{(2\pi)^{3/2}\sigma^{3}}e^{-\frac{\textbf{x}\cdot\textbf{x}}{2\sigma^{2}}}.$ (55)
This does not correspond to a physical quantized vortex, for which $\rho$ must vanish at the core; however, it is a simple and instructive example. In this case, the two confluent hypergeometric functions we need to evaluate are: $M(1,1,\kappa^{2})$ and $M(1,2,\kappa^{2})$. The first is simply given by:
$M(1,1,\kappa^{2})=e^{\kappa^{2}},$ (56)
which is easy to see from the definition of $M(a,b,z)$ in terms of the hypergeometric series:
$M(a,b,z)$ $=\sum_{k=0}^{\infty}\frac{(a)_{k}}{(b)_{k}}\frac{z^{k}}{k!},$
where $(a)_{k}$ and $(b)_{k}$ are Pochhammer symbols, where:
$(x)_{k}$ $=\frac{\Gamma(x+k)}{\Gamma(x)},$
so that:
$\frac{(a)_{k}}{(b)_{k}}$ $=\frac{\Gamma(a+k)}{\Gamma(a)}\frac{\Gamma(b)}{\Gamma(b+k)}.$
Note from this definition that $M(a,a,z)=e^{z}$. It is also straightforward to see:
$M(0,b,z)$ $=\sum_{k=0}^{\infty}\frac{\Gamma(k)}{\Gamma(0)}\frac{\Gamma(b)}{\Gamma(b+k)}\frac{z^{k}}{k!},$
$=1+\frac{1}{\Gamma(0)}\sum_{k=1}^{\infty}\frac{\Gamma(k)\Gamma(b)}{\Gamma(b+k)}\frac{z^{k}}{k!},$
$=1.$
With Eq. (56) and relation $(vi)$ from Eqs. (50), it is straightforward to find:
$M(1,2,\kappa^{2})=\frac{e^{\kappa^{2}}-1}{\kappa^{2}}.$ (57)
With Eqs. (56) and (57) we see that the coarse-grained variables for $n=0$ are:
$\left\langle\rho\right\rangle=e^{-\frac{z^{2}+R^{2}}{2\ell^{2}}\left(1-\frac{\tilde{\sigma}^{2}}{\ell^{2}}\right)}\frac{m\left(2\tilde{\sigma}^{2}\right)^{3/2}}{(4\pi)^{3/2}\ell^{3}\sigma^{3}},$ (58)
$\left\langle\left\langle\textbf{u}\right\rangle\right\rangle=\frac{\Gamma}{2\pi}\frac{1-e^{-\frac{R^{2}\tilde{\sigma}^{2}}{2\ell^{4}}}}{R\sigma^{2}}\left(\sigma^{2}+\ell^{2}\right)\ \hat{\theta}$, (59)
and
$\boldsymbol{\omega}=\hat{z}\frac{\Gamma}{2\pi\ell^{2}}e^{-\frac{R^{2}\tilde{\sigma}^{2}}{2\ell^{4}}}.$ (60)
Notice that, for $R=0$ the vorticity is $\boldsymbol{\omega}=\hat{z}\Gamma/2\pi\ell^{2}$ and, hence, diverges as $\ell\rightarrow 0$. This is expected due to the singular nature of vorticity in the microscopic problem.
A.2 Special Case: n=1
The case of $n=1$ corresponds to the density field:
$\rho=\frac{m\left(\textbf{x}_{\perp}\cdot\textbf{x}_{\perp}\right)}{(2\pi)^{3/2}2\sigma^{5}}e^{-\frac{\textbf{x}\cdot\textbf{x}}{2\sigma^{2}}}.$ (61)
This corresponds to a more physical representation of quantized vortex, for which $\rho\sim R^{2}$ near the vortex core. In this case, the two confluent hypergeometric functions we need to evaluate are: $M(2,1,\kappa^{2})$ and $M(2,2,\kappa^{2})$. The second of these is clearly:
$M(2,2,\kappa^{2})=e^{\kappa^{2}}.$ (62)
To find $M(2,1,\kappa^{2})$ we can again use relation $(vi)$ from Eqs. (50) to find:
$M(2,1,\kappa^{2})=(1+\kappa^{2})e^{\kappa^{2}}.$ (63)
With Eqs. (62) and (63) we see that the coarse-grained variables for $n=1$ are:
$\left\langle\rho\right\rangle=e^{-\frac{z^{2}+R^{2}}{2\ell^{2}}\left(1-\frac{\tilde{\sigma}^{2}}{\ell^{2}}\right)}\frac{m\left(2\tilde{\sigma}^{2}\right)^{5/2}}{(4\pi)^{3/2}\ell^{3}2\sigma^{5}}\left(1+\frac{R^{2}\sigma^{2}}{2\ell^{2}\left(\ell^{2}+\sigma^{2}\right)}\right),$ (64)
$\left\langle\left\langle\textbf{u}\right\rangle\right\rangle=\frac{R\ \Gamma}{2\pi}\frac{1}{2\ell^{2}+\tfrac{R^{2}\sigma^{2}}{\ell^{2}+\sigma^{2}}}\ \hat{\theta}$, (65)
and
$\boldsymbol{\omega}=\hat{z}\frac{\Gamma}{\pi}\frac{2\ell^{2}}{\left(2\ell^{2}+\tfrac{R^{2}\sigma^{2}}{\ell^{2}+\sigma^{2}}\right)^{2}}.$ (66)
At $R=0$, the vorticity is $\boldsymbol{\omega}=\hat{z}\Gamma/2\pi\ell^{2}$, just as the $n=0$ case. However, for $R,\sigma>0$, and $\ell<<R,\sigma$, the vorticity becomes:
$\boldsymbol{\omega}=\hat{z}\frac{2\Gamma}{\pi}\frac{\ell^{2}}{R^{4}}+\mathcal{O}(\ell^{4}).$ (67)
Comparing this to the finite scale expansion presented in the main text, we confirm that the leading-order contributions to the vorticity go as $\sim\ell^{2}$.